Suppose that there are only two goods in the economy, namely goods A and goods B. Use the 2014 as base year and use the information in the table below to calculate the real GDP, the GDP deflator, and the inflation rate by GDP deflator in 2015;

2014 2015
Price Quantity

Price

Quantity
A $31 10000 $36 1050
B $102 200 $100 205
Answer #1

Answer

real GDP =sum of (the price of the base year * quantity of the year)

real GDP in 2014=31*10000+102*200=330400

real GDP in 2015=31*1050+102*205=53460

===========

Nominal GDP =sum of (the price of the year *quantity of the year)

NGDP in 2014==31*10000+102*200=330400

NGDP in 2015=36*1050+100*205=58300

==========

GDP deflator =(NGDP/RGDP)*100

GDP deflator in 2014=(330400/330400)*100=100

GDP deflator in 2015=(58300/53460)*100=109.053498

Inflation =((GDP deflator of the year - GDP deflator of last year)/GDP deflator of last year)*100

=((109.053498-100)/100)*100

=9.053498

=9.05%